# The Racah algebra of rank 2: Properties, symmetries and representation

### v1

@@
Abstract
This paper is threefold. First, we provide a new “universal” definition for the Racah algebra of rank-2 as an extension of the rank-1 Racah algebra where the generators are indexed by subsets and any three disjoint indexing sets define a subalgebra isomorphic to the rank-1 case. With this definition, we explore some of the properties of the algebra including verifying that these natural assumptions are equivalent to other defining relations in the literature. Second, we look at the symmetries of the generators of the rank-2 Racah algebra. Those symmetries allows us to partially make abstraction of the choice of the generators and write relations and properties in a different format. Last, we provide a novel representation of the Racah algebra. This new representation requires only one generator to be diagonal and is based on an expansion of the split basis representation from the rank-1 Racah algebra.
@@ 1 Introduction
The Racah algebra and its generalization has recently been the subject of interest in representation theory and in connection with (multivariate) orthogonal polynomials, see for example the recent review article [8] and references therein. The title of the algebra refers to Guilio Racah, who defined what are now known as Racah coefficients to express the recoupling of triples of angular momentum [19]. The terminology is commonly used now to express recoupling coefficients of triples of various irreducible representations, including of quantum groups [16, 15]. Later the Racah polynomials, in particular their quantum versions, were defined by Askey and Wilson explicitly as ”A set of orthogonal polynomials that generalize the Racah coefficients or 6-j symbols”[1].
As is well known thanks to the work of Zhedanov, bispectral orthogonal polynomials admit quadratic algebras [9] and the Racah algebra is the algebra associated with the classical (q=1) Racah polynomials. Again, it was Zhedanov [10] who drew an explicit connection between the coefficients of Racah and the symmetry algebra. Now we understand the Racah algebra and the Racah polynomials in particular in terms of tensor products of $su(1,1)$ [12, 11, 18]. Due to the inductive nature of the tensor product, this construction naturally generalizes to higher rank [8, 7] and can be seen as the commutant of the diagonal embedding of $\mathfrak{sl}_{2}$ into the n-fold tensor product of its universal algebra[5].
From a different perspective, there has been work to define a “universal” version of the Racah algebra [2, 5], divorced from its realization in terms of tensor products, and to understand its representation theory [13]. This paper follows in this vein. In particular, we define the rank-2 Racah algebra $R(4)$ as an associative algebra with a set of generators and relations and show that it is equivalent to the definition given in [5]. In addition, we use the Jacobi identity to recover relation (3.9) of [5] and show that the Jacobi identity for all commutation relations are satisfied modulo this relation.
@@ 2.1 The rank-1 Racah algebra
Definition 1.
The Racah algebra of rank 1, denoted $R(3)$ is defined to be the unital, associative algebra over a field $\mathcal{K}$ with generators $C_{I}$ with $I\subset\{1,2,3\}$ satisfying the following relations
$[C_{i},C_{I}]=0\mbox{ for all }i=1,2,3,$ (2.1)
$[C_{123},C_{I}]=0,$ (2.2)
@@
$\frac{1}{2}[C_{jk},[C_{ij},C_{jk}]]=C_{ik}C_{jk}-C_{jk}C_{ij}+(C_{k}-C_{j})(C_{i}-C_{ijk}),\qquad i\neq j\neq k\neq i,$ (2.4)
where [,]normal-⋅normal-⋅[$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the regular commutator.
To simplify the notation, we omit the set notation and write $C_{ij}=C_{\{i,j\}}$, although it is important to remember that the generators $C_{I}$ are symmetric in the indices. Indeed, it is worth noting that the set of equations (2.12.4) are invariant under the group of permutation $\mathcal{P}_{3}$ acting on the indices. In addition, the maximal elements $C_{123}$ and the elements with one index, $C_{i}$, $i=1,2,3$ are in the center of the algebra. Furthermore, one should note that the Racah algebra is not a Lie algebra as it closes quadratically.
Often, an additional linearly independent element is introduced to avoid commutators of commutators, that is
$D_{123}\equiv\frac{1}{2}[C_{12},C_{23}],$ (2.5)
@@
$D_{123}=\frac{1}{2}[C_{23},C_{13}]=\frac{1}{2}[C_{13},C_{12}],$ (2.6)
as a consequence of relations (2.1-2.3). With this definition, we can express (2.4) as
@@
Proof: We need to check only the non-central terms $C_{12},$ $C_{23},$ $C_{13}$ and $C_{123}$. The first case to check is
@@
$=$ $2[C_{13},D_{123}]+2[C_{12},D_{123}]+2[C_{23},D_{123}]$
$=$ $2\times\left(C_{12}C_{13}-C_{13}C_{23}+(C_{1}-C_{3})(C_{2}-C_{123})\right.$
$\qquad+C_{23}C_{12}-C_{12}C_{13}+(C_{2}-C_{1})(C_{3}-C_{123})$
$\left.\qquad+C_{13}C_{23}-C_{23}C_{12}+(C_{3}-C_{2})(C_{1}-C_{123})\right)$
$=$ $0$
@@
The second (and last) case to check is a pair of $C$’s with the $D$. Up to permutation it is as follows
@@
We finish our discussion of the rank-1 case by noting that the algebra possesses a Casimir operator, that is
$\mathfrak{C}$ $=$ $D_{123}^{2}-\frac{1}{2}\{C_{12}^{2},C_{23}\}-\frac{1}{2}\{C_{23}^{2},C_{12}\}+C_{12}^{2}+C_{23}^{2}+\{C_{12},C_{23}\}$ (2.10)
$+\dfrac{1}{2}(C_{1}+C_{2}+C_{3}+C_{123})(\{C_{12},C_{23}]\}-2C_{12}-2C_{23})\}$
@@
$+(C_{1}+C_{3})(C_{123}+C_{2})+(C_{1}C_{3}-C_{123}C_{2})(C_{123}-C_{1}+C_{2}-C_{3}),$
where the last line involves only central elements. If one uses the $\mathfrak{su}(1,1)$ approach to construct a representation of a Racah algebra, see e.g. [12], then it has been shown that the Casimir will take a value of 0, which corresponds to the so-called special Racah. However, in this paper, we will not start from $\mathfrak{su}(1,1)$, but from a definition of the algebra, both for the rank-1 and rank-2 cases.
@@ 2.2 The rank-2 and higher Racah algebras
Definition 2.
The Racah algebra $R(n)$ is defined to be the unital associative algebra over a field $\mathcal{K}$ with generators $C_{I}$ with $I\subset\{1,2,3,\dots,{n}\}$ satisfying the following relations
$[C_{I},C_{J}]=0\mbox{ for all }I\subset J\mbox{ or }I\cap J=\emptyset,$ (2.11)
@@
$\frac{1}{2}[C_{JK},[C_{IJ},C_{JK}]]=C_{IK}C_{JK}-C_{JK}C_{IJ}+(C_{K}-C_{J})(C_{I}-C_{IJK}),$ (2.13)
where $I,J$ and $K$ are disjoint subsets in equations (2.12, 2.13). For notational convience, we omit the set notation and union sign and express $IJ=I\cup J.$
Lemma 2.
The following relations hold in the algebra $R(n)$
$[C_{IJ},C_{JK}]=[C_{KI},C_{IJ}]=[C_{JK},C_{KI}].$ (2.14)
Proof This follows directly from the commutation (2.11) and decomposition relations (2.12). ∎
Lemma 3.
The following relation holds in the algebra $R(4)$
$\frac{1}{2}[C_{KI},[C_{IJ},C_{JK}]]=C_{IJ}C_{KI}-C_{KI}C_{JK}+(C_{I}-C_{K})(C_{J}-C_{IJK}).$ (2.15)
Proof: From (2.13) with $K\rightarrow I\rightarrow J$ we have
$\frac{1}{2}[C_{KI},[C_{JK},C_{KI}]]=C_{IJ}C_{KI}-C_{KI}C_{JK}+(C_{I}-C_{K})(C_{J}-C_{IJK}).$
The resulting identity follows from the relations in Lemma 2. $\square$
Lemma 4.
@@
Proof: This proof is directly analogous to Lemma 1 with indices replaced by subset. Relation (2.13) is a straightforward computation that follows from symmetry of the commutator
@@
$=$ $2\times\left(C_{IJ}C_{KI}-C_{KI}C_{JK}+(C_{I}-C_{K})(C_{J}-C_{IJK})\right.$
$\qquad+C_{JK}C_{IJ}-C_{IJ}C_{KI}+(C_{J}-C_{I})(C_{K}-C_{IJK})$
$\left.\qquad+C_{KI}C_{JK}-C_{JK}C_{IJ}+(C_{K}-C_{J})(C_{I}-C_{IJK})\right)$
$=$ $0$
Thus the triple commutation of any of the generators satisfies the Jacobi identity identically, as expected in an associative algebra. $\square$
@@
$D_{ijk}\equiv\frac{1}{2}[C_{ij},C_{jk}],\qquad 1\leq i<j<k\leq n.$
Note that unlike the $C$’s, these elements are defined for a trio of distinct ordered indices. However, we see that they are invariant under cyclic permutations as a result of Lemma 2:
$D_{ijk}=\frac{1}{2}[C_{jk},C_{ik}]=\frac{1}{2}[C_{ik},C_{ij}].$
@@
Lemma 5.
The following commutation relations hold in $R(n)$
$[C_{k\ell},D_{ijk}]+[C_{k\ell},D_{ij\ell}]=0.$ (2.18)
Proof. Note that $C_{k\ell}$ and hence $P_{k\ell}$ will commute with both $C_{ik\ell}$ and $C_{jk\ell}.$ We use the decomposition relation (2.12) to obtain
@@
$=2\left(D_{ikj}+D_{ik\ell}+D_{\ell kj}+D_{k\ell j}+D_{i\ell k}+D_{i\ell j}\right).$
Next we use the fact that $D_{ijk}$ is invariant under cyclic permutations and odd under singleton flips $D_{ijk}=-D_{ikj}.$ This leaves
$[C_{ik\ell},C_{jk\ell}]=-2D_{ijk}-2D_{ij\ell}.$ (2.19)
So, finally we see that Lemma 5 holds from the commutativity of $C_{k\ell}$ with $C_{ik\ell}$ and $C_{jk\ell}.$
Of course, this lemma does not give us the required relation $[C_{k\ell},D_{ijk}]$ rather relations between such terms. For example in $R(4)$, we are looking to obtain 24 equation with leading order terms $[P_{A},D_{B}]$ with $A$ a two element subset and $B$ a three element subset. We have 12 “inner” equations of the form (2.17). There are 6 equations coming from Lemma 5 plus 4 from the fact that $C_{ijk\ell}$ is central and 4 from the fact that $C_{ijk}$ commutes with $D_{ijk}.$ Finally, there are 12 additional choices of $I,J,K$ where at least one subset contains 3 elements. This leads to consistent linear system that can be solved for the required relations.
Lemma 6.
In the Racah algebra $R(n)$, the elements $P_{ij}$ and $D_{jk\ell}$ satisfy the following relation
@@
Proof. As described above, we used a symbolic manipulator (MAPLE [17]) to solve the linear system of equations involving terms $P_{A}D_{B}$ for the required expressions. The resulting identities are non-trivial but can be checked as follows. On the one hand, it can be directly verified using the commutation (2.11) and decomposition relations (2.12) that $[C_{ij},D_{ijk}]$ can be expressed as
$[C_{ij},D_{jkl}]$ $=$ $-4[C_{ijk\ell}-C_{ik\ell},D_{ik\ell}]+2[C_{ijk\ell}-C_{jk\ell},D_{jk\ell}]$
$+2[C_{ijk\ell}-C_{ijk},D_{ijk}]-2[C_{ijk\ell}-C_{ij\ell},D_{ij\ell}]$
$+4[C_{ij},D_{jk\ell}+D_{ik\ell}]-2[C_{ik},D_{jkl}-D_{ij\ell}]-2[C_{i\ell},D_{jk\ell}+D_{ijk}]$
$-4[C_{k\ell},D_{ik\ell}]-2[C_{j\ell},D_{ij\ell}]+2[C_{jk},D_{ijk}]$
@@
On the other hand, the first 2 lines are equal to 0 as consequences of the commutation relations (2.11), as $C_{ijkl}$ and $C_{ijk}$ will both commute with $D_{ijk}$ (and permutations). The third line also vanishes from the relation in Lemmma 5. Finally, the second to last line is comprised of terms containing “inner” quadratic relations for singleton subsets (2.17) and the last line “inner” quadratic relations for multindex sets $I=\{i\},J=\{j\},K=\{k\ell\}$ and $I=\{i\},J=\{k\},K=\{j\ell\}$ respectively in (2.13). Replacing these last two lines with their quadratic identities leaves the resulting equation
$[C_{ij},D_{jk\ell}]=(C_{i\ell}-C_{i}-C_{j})(C_{jk}-C_{j}-C_{k})-(C_{j\ell}-C_{j}-C_{\ell})(C_{ik}-C_{i}-C_{k})$
@@
Lemma 7.
The outer commutators satisfy the following identities
$[P_{ij},D_{jk\ell}]=-[P_{ji},D_{ik\ell}]$ (2.22)
$[P_{ij},D_{jk\ell}]=[P_{k\ell},D_{\ell ij}]$ (2.23)
@@
Proof: Each of these are direct consequences of Lemma 6. We also note that the first equation is exactly Lemma 5 expressed in terms of $P$. ∎
For the final relations we would like to understand the commutation of the $D$’s with each other. These can be determined from the definition of the $D$’s plus the assumption that we have an associative algebra and so the Jacobi relations hold. In particular we have the following identity.
Lemma 8.
In the Racah algebra $R(n)$, the following relation holds
@@
Proof: We use Lemma 2 to write
$[D_{ijk},D_{jk\ell}]=\frac{1}{2}[D_{ijk},[P_{k\ell},P_{\ell j}]]$
using the fact that our algebra is associative and so the Jacobi identity holds gives
$[D_{ijk},D_{jk\ell}]$ $=$ $\frac{1}{2}[P_{\ell j},[P_{k\ell},D_{ijk}]]-\frac{1}{2}[P_{k\ell},[P_{\ell j},D_{ijk}]]$
$=$ $\frac{1}{2}[P_{\ell j},[P_{\ell k},D_{kij}]]-\frac{1}{2}[P_{k\ell},[P_{\ell j},D_{jki}]]$
$=$ $\frac{1}{2}[P_{\ell j},P_{\ell j}P_{ki}-P_{kj}P_{\ell i}]-\frac{1}{2}[P_{k\ell},P_{\ell i}P_{jk}-P_{ji}P_{\ell k}]$
$=$ $-\frac{1}{2}[P_{\ell j},P_{jk}P_{\ell i}]-\frac{1}{2}[P_{k\ell},P_{jk}P_{\ell i}]$
$=$ $-D_{\ell jk}P_{\ell i}-P_{jk}D_{j\ell i}-D_{\ell kj}P_{\ell i}-P_{jk}D_{k\ell i}$
$=$ $-P_{jk}(D_{j\ell i}+D_{k\ell i}).$
@@
$[D_{ijk},D_{jk\ell}]=P_{jk}\left(D_{ji\ell}+D_{ki\ell}\right).$ (2.26)
Note that we can also apply Lemma 5 to reverse the order of the right-hand side, giving the alternate form
@@
Of course, there are two other expressions that we could have used for the definition of the $D_{jk\ell}.$ The fact that each of these are equal is equivalent to verifying the Jacobi identity triples with triples of the form $[D,[P,P]]$. We will return to problem in more generality but for now we consider the following
@@
$-[P_{k\ell},[P_{jk},D_{ijk}]]=-[P_{k\ell},\left(P_{jk}-2P_{j}\right)P_{ik}-P_{ij}\left(P_{jk}-2P_{k}\right)]$
$=-2D_{\ell kj}P_{ik}-2\left(P_{jk}+2P_{j}\right)D_{\ell ki}+2P_{ij}D_{\ell kj}$
@@
Finally, we note that we can use (2.24) from Lemma 7 to express the relation with the order switched,
$0=2D_{jk\ell}P_{i}+D_{ik\ell}P_{ji}+D_{i\ell j}P_{ki}+D_{ijk}P_{\ell i}.$ (2.30)
Thus we see that this identity can be understood as a consequence of the Jacobi identity for the commutations relations for the generators, as was first shown [5]. We record the previous results in the following theorem and show that the remaining Jacobi relations are satisfied without additional constraints in the rank-2 case, $R(4)$. In particular, we mention that Theorem 1 shows that our definition of the Racah algebra is equivalent to the definition given in [5].
Theorem 1.
The Racah algebra $R(4)$ is generated by the elements $P_{i}$, $P_{ij}$ (2.16) and $D_{ijk}$ where the $P$’s are symmetric under interchange of the indices and the $D$’s are anti-symmetric. The generators satisfy the following commutation relations

### v2

@@
[…]
The Racah Algebra of Rank 2: Properties, Symmetries and Representation
[…]
The Racah Algebra of Rank 2:
Properties, Symmetries and Representation
[…]
Received March 08, 2024, in final form September 10, 2024; Published online September 22, 2024
[…]
The goals of this paper are threefold. First, we provide a new “universal” definition for the Racah algebra of rank 2 as an extension of the rank-1 Racah algebra where the generators are indexed by subsets and any three disjoint indexing sets define a subalgebra isomorphic to the rank-1 case. With this definition, we explore some of the properties of the algebra including verifying that these natural assumptions are equivalent to other defining relations in the literature. Second, we look at the symmetries of the generators of the rank-2 Racah algebra. Those symmetries allows us to partially make abstraction of the choice of the generators and write relations and properties in a different format. Last, we provide a novel representation of the Racah algebra. This new representation requires only one generator to be diagonal and is based on an expansion of the split basis representation from the rank-1 Racah algebra.
[…]
Racah algebra; representation; symmetry; rank-2
[…]
81R12; 20C35
@@ 1 Introduction
The Racah algebra and its generalization has recently been the subject of interest in representation theory and in connection with (multivariate) orthogonal polynomials, see for example the recent review article [6] and references therein. The title of the algebra refers to Giulio Racah, who defined what are now known as Racah coefficients to express the recoupling of triples of angular momentum [18]. The terminology is commonly used now to express recoupling coefficients of triples of various irreducible representations, including of quantum groups [13, 14, 15]. Back in the Istanbul Summer School [19], Racah even proposed an algebra that was closing of degree 3, compared to the Racah algebra that closes quadratically. Later the Racah polynomials, in particular their quantum versions, were defined by Askey and Wilson explicitly as “A set of orthogonal polynomials that generalize the Racah coefficients or $6$-j symbols” [1].
As is well known thanks to the work of Zhedanov, bispectral orthogonal polynomials admit quadratic algebras [9] and the Racah algebra is the algebra associated with the classical ${(q=1)}$ Racah polynomials. Again, it was Zhedanov [8] who drew an explicit connection between the coefficients of Racah and the symmetry algebra. Now we understand the Racah algebra and the Racah polynomials in particular in terms of tensor products of $\mathfrak{sl}_{2}(\mathbb{C})$ and its real form $\mathfrak{su}(1,1)$ [7, 10, 17]. Due to the inductive nature of the tensor product, this construction naturally generalizes to higher rank [5, 6] and can be seen as the commutant of the diagonal embedding of $\mathfrak{sl}_{2}(\mathbb{C})$ into the $n$-fold tensor product of its universal algebra [4].
From a different perspective, there has been work to define a “universal” version of the Racah algebra [2, 4], divorced from its realization in terms of tensor products, and to understand its representation theory [11]. This paper follows in this vein. In particular, we define the rank-2 Racah algebra $R(4)$ as an associative algebra with a set of generators and relations and show that it is equivalent to the definition given in [4]. In addition, we use the Jacobi identity to recover relation (3.9) of [4] and show that the Jacobi identity for all commutation relations is satisfied modulo this relation.
@@ 2.1 The rank-1 Racah algebra
Definition 2.1.
The Racah algebra of rank $1$, denoted $R(3)$ is defined to be the unital, associative algebra over a field $\mathcal{K}$ with generators $C_{I}$ with $I\subset\{1,2,3\}$ satisfying the following relations:
$[C_{i},C_{I}]=0\quad\mbox{for all}\quad i=1,2,3,\qquad[C_{123},C_{I}]=0,$ (2.1)
@@
$\tfrac{1}{2}[C_{jk},[C_{ij},C_{jk}]]=C_{ik}C_{jk}-C_{jk}C_{ij}+(C_{k}-C_{j})(C_{i}-C_{ijk}),\qquad i\neq j\neq k\neq i,$ (2.3)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the regular commutator.
To simplify the notation, we omit the set notation and write $C_{ij}=C_{\{i,j\}}$, although it is important to remember that the generators $C_{I}$ are symmetric in the indices. Indeed, it is worth noting that the set of equations (2.1)–(2.3) are invariant under the group of permutation $\mathcal{P}_{3}$ acting on the indices. In addition, the maximal elements $C_{123}$ and the elements with one index, $C_{i}$, $i=1,2,3$, are in the center of the algebra. Furthermore, one should note that the Racah algebra is not a Lie algebra as it closes quadratically.
Often, an additional linearly independent element is introduced to avoid commutators of commutators, that is,
$D_{123}\equiv\tfrac{1}{2}[C_{12},C_{23}],$ (2.4)
@@
$D_{123}=\tfrac{1}{2}[C_{23},C_{13}]=\tfrac{1}{2}[C_{13},C_{12}],$ (2.5)
as a consequence of relations (2.1)–(2.2). With this definition, we can express (2.3) as
@@
Proof.
We need to check only the non-central terms $C_{12}$, $C_{23}$, $C_{13}$ and $D_{123}$. The first case to check is
@@
$\qquad=2[C_{13},D_{123}]+2[C_{12},D_{123}]+2[C_{23},D_{123}]$
$\qquad=2\left(C_{12}C_{13}-C_{13}C_{23}+(C_{1}-C_{3})(C_{2}-C_{123})+C_{23}C_{12}-C_{12}C_{13}\right.$
$\phantom{=}{}\qquad\left.+(C_{2}-C_{1})(C_{3}-C_{123})+C_{13}C_{23}-C_{23}C_{12}+(C_{3}-C_{2})(C_{1}-C_{123})\right)=0.$
@@
The second (and last) case to check is a pair of $C$’s with the $D$. Up to permutation, it is as follows:
@@
We finish our discussion of the rank-1 case by noting that the algebra possesses a Casimir operator, that is,
$\mathfrak{C}=D_{123}^{2}-\tfrac{1}{2}\big{\{}C_{12}^{2},C_{23}\big{\}}-\tfrac{1}{2}\big{\{}C_{23}^{2},C_{12}\big{\}}+C_{12}^{2}+C_{23}^{2}+\{C_{12},C_{23}\}$
${}+\tfrac{1}{2}(C_{1}+C_{2}+C_{3}+C_{123})(\{C_{12},C_{23}\}-2C_{12}-2C_{23})$
@@
${}+(C_{1}+C_{3})(C_{123}+C_{2})+(C_{1}C_{3}-C_{123}C_{2})(C_{123}-C_{1}+C_{2}-C_{3}),$ (2.7)
where the last line involves only central elements. If one uses the $\mathfrak{su}(1,1)$ approach to construct a representation of a Racah algebra, see, e.g., [7], then it has been shown that the Casimir operator will take a value of 0, which corresponds to the so-called special Racah. However, in this paper, we will not start from $\mathfrak{su}(1,1)$, but from a definition of the algebra, both for the rank-1 and rank-2 cases.
@@ 2.2 The rank-2 and higher Racah algebras
Definition 2.3.
The Racah algebra $R(n)$ is defined to be the unital associative algebra over a field $\mathcal{K}$ with generators $C_{I}$ with $I\subset\{1,2,3,\dots,{n}\}$ satisfying the following relations:
$[C_{I},C_{J}]=0\quad\mbox{for all}\quad I\subset J\quad\mbox{or}\quad I\cap J=\varnothing,$ (2.8)
@@
$\tfrac{1}{2}[C_{JK},[C_{IJ},C_{JK}]]=C_{IK}C_{JK}-C_{JK}C_{IJ}+(C_{K}-C_{J})(C_{I}-C_{IJK}),$ (2.10)
where $I$, $J$ and $K$ are disjoint subsets in equations (2.9), (2.10). For notational convenience, we omit the set notation and union sign and express $IJ=I\cup J$.
Lemma 2.4.
The following relations hold in the algebra $R(n)$:
$[C_{IJ},C_{JK}]=[C_{KI},C_{IJ}]=[C_{JK},C_{KI}].$
Proof.
This follows directly from the commutation (2.8) and decomposition relations (2.9). ∎
Lemma 2.5.
The following relation holds in the algebra $R(4)$:
$\tfrac{1}{2}[C_{KI},[C_{IJ},C_{JK}]]=C_{IJ}C_{KI}-C_{KI}C_{JK}+(C_{I}-C_{K})(C_{J}-C_{IJK}).$ (2.11)
Proof.
From (2.10) with $K\rightarrow I\rightarrow J$, we have
$\tfrac{1}{2}[C_{KI},[C_{JK},C_{KI}]]=C_{IJ}C_{KI}-C_{KI}C_{JK}+(C_{I}-C_{K})(C_{J}-C_{IJK}).$
The resulting identity follows from the relations in Lemma 2.4. ∎
Lemma 2.6.
@@
Proof.
This proof is directly analogous to Lemma 2.2 with indices replaced by subset. Relation (2.10) is a straightforward computation that follows from symmetry of the commutator
@@
$\qquad=2(C_{IJ}C_{KI}-C_{KI}C_{JK}+(C_{I}-C_{K})(C_{J}-C_{IJK})+C_{JK}C_{IJ}-C_{IJ}C_{KI}$
$\phantom{\qquad=}{}+(C_{J}-C_{I})(C_{K}-C_{IJK})+C_{KI}C_{JK}-C_{JK}C_{IJ}+(C_{K}-C_{J})(C_{I}-C_{IJK}))=0.$
Thus the triple commutation of any of the generators satisfies the Jacobi identity identically, as expected in an associative algebra. ∎
@@
$D_{ijk}\equiv\tfrac{1}{2}[C_{ij},C_{jk}],\qquad 1\leq i<j<k\leq n.$
Note that unlike the $C$’s, these elements are defined for a trio of distinct ordered indices. However, we see that they are invariant under cyclic permutations as a result of Lemma 2.4
$D_{ijk}=\tfrac{1}{2}[C_{jk},C_{ik}]=\tfrac{1}{2}[C_{ik},C_{ij}].$
@@
Lemma 2.7.
The following commutation relations hold in $R(n)$:
$[C_{k\ell},D_{ijk}]+[C_{k\ell},D_{ij\ell}]=0.$
Proof.
Note that $C_{k\ell}$ and hence $P_{k\ell}$ will commute with both $C_{ik\ell}$ and $C_{jk\ell}$. We use the decomposition relation (2.9) to obtain
@@
$=2(D_{ikj}+D_{ik\ell}+D_{\ell kj}+D_{k\ell j}+D_{i\ell k}+D_{i\ell j}).$
Next we use the fact that $D_{ijk}$ is invariant under cyclic permutations and odd under singleton flips $D_{ijk}=-D_{ikj}$. This leaves $[C_{ik\ell},C_{jk\ell}]=-2D_{ijk}-2D_{ij\ell}$. So, finally we see that Lemma 2.7 holds from the commutativity of $C_{k\ell}$ with $C_{ik\ell}$ and $C_{jk\ell}$. ∎
Of course, this lemma does not give us the required relation $[C_{k\ell},D_{ijk}]$ rather relations between such terms. For example in $R(4)$, we are looking to obtain 24 equation with leading order terms $[P_{A},D_{B}]$ with $A$ a two element subset and $B$ a three element subset. We have 12 “inner” equations of the form (2.13). There are 6 equations coming from Lemma 2.7 plus 4 from the fact that $C_{ijk\ell}$ is central and 4 from the fact that $C_{ijk}$ commutes with $D_{ijk}$. Finally, there are 12 additional choices of $I$, $J$, $K$ where at least one subset contains 3 elements. This leads to consistent linear system that can be solved for the required relations.
Lemma 2.8.
In the Racah algebra $R(n)$, the elements $P_{ij}$ and $D_{jk\ell}$ satisfy the following relation:
@@
Proof.
As described above, we used a symbolic manipulator (Maple [16]) to solve the linear system of equations involving terms $P_{A}D_{B}$ for the required expressions. The resulting identities are non-trivial but can be checked as follows. On the one hand, it can be directly verified using the commutation (2.8) and decomposition relations (2.9) that $[C_{ij},D_{ijk}]$ can be expressed as
$[C_{ij},D_{jkl}]=-4[C_{ijk\ell}-C_{ik\ell},D_{ik\ell}]+2[C_{ijk\ell}-C_{jk\ell},D_{jk\ell}]+2[C_{ijk\ell}-C_{ijk},D_{ijk}]$
${}-2[C_{ijk\ell}-C_{ij\ell},D_{ij\ell}]+4[C_{ij},D_{jk\ell}+D_{ik\ell}]-2[C_{ik},D_{jkl}-D_{ij\ell}]$
${}-2[C_{i\ell},D_{jk\ell}+D_{ijk}]-4[C_{k\ell},D_{ik\ell}]-2[C_{j\ell},D_{ij\ell}]+2[C_{jk},D_{ijk}]$
@@
On the other hand, the first 2 lines are equal to 0 as consequences of the commutation relations (2.8), as $C_{ijkl}$ and $C_{ijk}$ will both commute with $D_{ijk}$ (and permutations). The third line also vanishes from the relation in Lemma 2.7. Finally, the second to last line is comprised of terms containing “inner” quadratic relations for singleton subsets (2.13) and the last line “inner” quadratic relations for multi-index sets $I=\{i\}$, $J=\{j\}$, $K=\{k\ell\}$ and $I=\{i\}$, $J=\{k\}$, $K=\{j\ell\}$ respectively in (2.10). Replacing these last two lines with their quadratic identities leaves the resulting equation
$[C_{ij},D_{jk\ell}]=(C_{i\ell}-C_{i}-C_{j})(C_{jk}-C_{j}-C_{k})-(C_{j\ell}-C_{j}-C_{\ell})(C_{ik}-C_{i}-C_{k}),$
@@
Lemma 2.9.
The outer commutators satisfy the following identities:
$[P_{ij},D_{jk\ell}]=-[P_{ji},D_{ik\ell}],\qquad[P_{ij},D_{jk\ell}]=[P_{k\ell},D_{\ell ij}]$
@@
Proof.
Each of these are direct consequences of Lemma 2.8. We also note that the first equation is exactly Lemma 2.7 expressed in terms of $P$. ∎
For the final relations, we would like to understand the commutation of the $D$’s with each other. These can be determined from the definition of the $D$’s plus the assumption that we have an associative algebra and so the Jacobi relations hold. In particular, we have the following identity.
Lemma 2.10.
In the Racah algebra $R(n)$, the following relation holds:
@@
Proof.
We use Lemma 2.4 to write $[D_{ijk},D_{jk\ell}]=\frac{1}{2}[D_{ijk},[P_{k\ell},P_{\ell j}]]$ using the fact that our algebra is associative and so the Jacobi identity holds gives
$[D_{ijk},D_{jk\ell}]$ $=\tfrac{1}{2}[P_{\ell j},[P_{k\ell},D_{ijk}]]-\tfrac{1}{2}[P_{k\ell},[P_{\ell j},D_{ijk}]]$
$=\tfrac{1}{2}[P_{\ell j},[P_{\ell k},D_{kij}]]-\tfrac{1}{2}[P_{k\ell},[P_{\ell j},D_{jki}]]$
$=\tfrac{1}{2}[P_{\ell j},P_{\ell j}P_{ki}-P_{kj}P_{\ell i}]-\tfrac{1}{2}[P_{k\ell},P_{\ell i}P_{jk}-P_{ji}P_{\ell k}]$
$=-\tfrac{1}{2}[P_{\ell j},P_{jk}P_{\ell i}]-\frac{1}{2}[P_{k\ell},P_{jk}P_{\ell i}]$
$=-D_{\ell jk}P_{\ell i}-P_{jk}D_{j\ell i}-D_{\ell kj}P_{\ell i}-P_{jk}D_{k\ell i}=-P_{jk}(D_{j\ell i}+D_{k\ell i}).$
@@
$[D_{ijk},D_{jk\ell}]=P_{jk}\left(D_{ji\ell}+D_{ki\ell}\right).$
Note that we can also apply Lemma 2.7 to reverse the order of the right-hand side, giving the alternate form
@@
Of course, there are two other expressions that we could have used for the definition of the $D_{jk\ell}$. The fact that each of these is equal is equivalent to verifying the Jacobi identity triples with triples of the form $[D,[P,P]]$. We will return to problem in more generality but for now we consider the following:
@@
$-[P_{k\ell},[P_{jk},D_{ijk}]]$ $=-[P_{k\ell},(P_{jk}-2P_{j})P_{ik}-P_{ij}(P_{jk}-2P_{k})]$
$=-2D_{\ell kj}P_{ik}-2(P_{jk}+2P_{j})D_{\ell ki}+2P_{ij}D_{\ell kj}.$
@@
Finally, we note that we can use (2.15) from Lemma 2.9 to express the relation with the order switched,
$0=2D_{jk\ell}P_{i}+D_{ik\ell}P_{ji}+D_{i\ell j}P_{ki}+D_{ijk}P_{\ell i}.$
Thus we see that this identity can be understood as a consequence of the Jacobi identity for the commutations relations for the generators, as was first shown in [4]. We record the previous results in the following theorem and show that the remaining Jacobi relations are satisfied without additional constraints in the rank-2 case, $R(4)$. In particular, we mention that Theorem 2.11 shows that our definition of the Racah algebra is equivalent to the definition given in [4].
Theorem 2.11.
The Racah algebra $R(4)$ is generated by the elements $P_{i}$, $P_{ij}$ (2.12) and $D_{ijk}$ where the $P$’s are symmetric under interchange of the indices and the $D$’s are anti-symmetric. The generators satisfy the following commutation relations: